If a bond is worth $2490.00 what is it worth if it decreases by 33-1/3%
Answer by stanbon(75887)

If a bond is worth $2490.00 what is it worth if it decreases by 33-1/3%
-------
Note:: 1 - (33 1/3)% = 66 2/3%
------
Ans: (66.6666..)%*2490 = 0.66666666*2490 = 1660
-----------------------
Cheers,
Stan H.
-----------------